x^2-2x+2 devided by x+2

1
Expert's answer
2021-06-23T14:15:10-0400
"x^2-2x+2=x^2+2x-4x-8+10"

"=x(x+2)-4(x+2)+10"

"\dfrac{x^2-2x-2}{b}=\dfrac{x(x+2)-4(x+2)+10}{x+2}"

"=x-4+\dfrac{10}{x+2}, x\not=-2"